# Faster Diffusion Models via Higher-Order Approximation

### v1

@@ B.3 Proof of Lemma 4
Moreover, recognizing that for all $|\tau^{\prime}-\widetilde{\tau}|\leq c_{0}\widetilde{\tau}(1-\widetilde{\tau})$,
@@
provided that $|\tau^{\prime}-\widetilde{\tau}|\leq c_{0}\widetilde{\tau}(1-\widetilde{\tau})$. Here, the second inequality makes use of $\log(1+x)\leq x$. Recalling the definition of $\theta^{\prime}$, we know that
@@
Therefore, for all $|\tau^{\prime}-\widetilde{\tau}|\leq c_{0}\widetilde{\tau}(1-\widetilde{\tau})$, we have
@@ B.4 Proof of Lemma 6
=$\quad\stackrel{{\scriptstyle v=\frac{\sqrt{1-\overline{\alpha}_{t-1}}\log w}{\|u\|_{2}}}}{{=}}1+\int_{0}^{\infty}e^{\frac{\|u\|_{2}v}{\sqrt{1-\overline{\alpha}_{t-1}}}}\frac{\|u\|_{2}}{\sqrt{1-\overline{\alpha}_{t-1}}}\mathbb{P}\left(\|x_{t-1}^{\star}-\sqrt{\overline{\alpha}_{t-1}}X_{0}\|_{2}>\sqrt{1-\overline{\alpha}_{t-1}}v~{}\Big{|}~{}X_{t-1}=x_{t-1}^{\star}\right){\rm d}v$
@@
=$\stackrel{{\scriptstyle x_{\tau_{t},0}=x_{\tau_{t},0}^{\star}}}{{=}}\frac{p_{\frac{1}{\sqrt{1-\tau_{t,K-1}}}Y_{t-1}}\big{(}\frac{1}{\sqrt{1-\tau_{t,K-1}}}x_{\tau_{t,K-1}^{(N)}}\big{)}/p_{\frac{1}{\sqrt{1-\tau_{t,0}}}Y_{t}}\big{(}\frac{1}{1-\tau_{t,0}}x_{\tau_{t,0}}\big{)}}{p_{\frac{1}{\sqrt{\overline{\alpha}_{t-1}}}Y_{t-1}^{\star}}\big{(}\frac{1}{\sqrt{\overline{\alpha}_{t-1}}}x_{t-1}^{\star}\big{)}/p_{\frac{1}{\sqrt{1-\tau_{t,0}}}Y_{t}}\big{(}\frac{1}{1-\tau_{t,0}}x_{\tau_{t,0}}^{\star}\big{)}}$
@@ B.5 Proof of Lemma 8
We prove Lemma 8 by induction. For notational convenience, we let $\theta_{t}:=\theta_{\tau_{t,0}}(x_{\tau_{t,i}}^{(0)})$. Recalling that $x_{\tau}^{\star}$ is the solution of ODE (9) at $\tau$ with the initial condition $x_{\tau_{t,0}}^{\star}=x_{\tau_{t,0}}$, we know from Lemma 4 that $-\log p_{\overline{X}_{\tau}}(x_{\tau}^{\star})\leq 2\theta_{t}d\log T$ for all $\tau_{t,K-1}\leq\tau\leq\tau_{t,0}$. For all $\lambda\in[0,1]$, one can derive
@@
[…]
@@
=(8)${{=}}1-\frac{\tau_{t,0}-\tau_{t,i}}{2(1-\tau_{t,0})\tau_{t,0}\tau_{t,i}}\bigg{[}\int_{x_{0}}p_{X_{0}\,|\,\overline{X}_{\tau_{t,0}}}(x_{0}\,|\,x_{\tau_{t,0}})\left\|x_{\tau_{t,0}}-\sqrt{1-\tau_{t,0}}x_{0}\right\|_{2}^{2}{\rm d}x_{0}$
@@
Repeating similar arguments and using induction yields that for all $10\leq n\leq N-$ 1, $10\leq i\leq K-$ 1,
@@ B.6 Proof of Lemma 9
We prove Lemma 9 by contradiction. Suppose that there exists $\ell\in[1,\tau(x_{T}))$ such that $-\log q_{\ell}(x_{\ell})>2c_{6}d\log T$. We let $1<t\leq\ell$ denote the smallest time step satisfying

### v2

@@ B.3 Proof of Lemma 4
[…]
@@
[…]
@@
[…]
@@ B.4 Proof of Lemma 6
$\quad\stackrel{{\scriptstyle v=\frac{\sqrt{1-\overline{\alpha}_{t-1}}\log w}{\|u\|_{2}}}}{{=}}1+\int_{0}^{\infty}e^{\frac{\|u\|_{2}v}{\sqrt{1-\overline{\alpha}_{t-1}}}}\frac{\|u\|_{2}}{\sqrt{1-\overline{\alpha}_{t-1}}}\mathbb{P}\left(\|x_{t-1}^{\star}-\sqrt{\overline{\alpha}_{t-1}}X_{0}\|_{2}>\sqrt{1-\overline{\alpha}_{t-1}}v~\Big{|}~X_{t-1}=x_{t-1}^{\star}\right){\rm d}v$
@@
$\stackrel{{\scriptstyle x_{\tau_{t},0}=x_{\tau_{t},0}^{\star}}}{{=}}\frac{p_{\frac{1}{\sqrt{1-\tau_{t,K-1}}}Y_{t-1}}\big{(}\frac{1}{\sqrt{1-\tau_{t,K-1}}}x_{\tau_{t,K-1}^{(N)}}\big{)}/p_{\frac{1}{\sqrt{1-\tau_{t,0}}}Y_{t}}\big{(}\frac{1}{1-\tau_{t,0}}x_{\tau_{t,0}}\big{)}}{p_{\frac{1}{\sqrt{\overline{\alpha}_{t-1}}}Y_{t-1}^{\star}}\big{(}\frac{1}{\sqrt{\overline{\alpha}_{t-1}}}x_{t-1}^{\star}\big{)}/p_{\frac{1}{\sqrt{1-\tau_{t,0}}}Y_{t}}\big{(}\frac{1}{1-\tau_{t,0}}x_{\tau_{t,0}}^{\star}\big{)}}$
@@ B.5 Proof of Lemma 8
We prove Lemma 8 by induction. For notational convenience, we let $\theta_{t}:=\theta_{\tau_{t,0}}(x_{\tau_{t,i}}^{(0)})$. Recalling that $x_{\tau}^{\star}$ is the solution of ODE (9) at $\tau$ with the initial condition $x_{\tau_{t,0}}^{\star}=x_{\tau_{t,0}}$, we know from Lemma 4 that $\log p_{\overline{X}_{\tau}}(x_{\tau}^{\star})\leq 2\theta_{t}d\log T$ for all $\tau_{t,K-1}\leq\tau\leq\tau_{t,0}$. For all $\lambda\in[0,1]$, one can derive
@@
$\bigg{\|}-\sqrt{1-\tau_{t,0}}\int_{\tau_{t,0}}^{\tau_{t,i}}\frac{1}{2(1-\tau)^{3/2}}s_{\tau}^{\star}(x_{\tau}^{\star})\mathrm{d}\tau\bigg{\|}_{2}$ $\lesssim\sqrt{1-\tau_{t,0}}\int_{\tau_{t,i}}^{\tau_{t,0}}\frac{1}{2(1-\tau)^{3/2}}\sqrt{\frac{d\theta_{t}\log T}{\tau}}d\tau$
@@
[…]
@@
Repeating similar arguments and using induction yields that for all $0\leq n\leq N-1$, $0\leq i\leq K-1$,
@@ B.6 Proof of Lemma 9
We prove Lemma 9 by contradiction. Suppose that there exists $\ell\in[1,\tau(x_{T}))$ such that $\log q_{\ell}(x_{\ell})>2c_{6}d\log T$. We let $1<t\leq\ell$ denote the smallest time step satisfying